an airplane travels 400 miles against the wind in 4 hours, and makes the return trip with the same wind in 1 hours. find the rate of the wind
Answer by nyc_function(2741)

Let x = speed of airplane
Let y = speed of wind
Against:
rate = x - y
time = 4
distance = 400
4(x - y) = 400
4x - 4y = 400....Thisis your first equation.
============================================
With Wind:
rate = x + y
time = 1
distance = 400
x + y = 400....This is your second equation.
============================================
Here they are together:
4x - 4y = 400
x + y = 400
Can you finish now?